### v1

# Enhancing Memory and Imagination Consistency in Diffusion-based World Models via Linear-Time Sequence Modeling

Abstract
World models are crucial for enabling agents to simulate and plan within environments, yet existing approaches struggle with long-term dependencies and inconsistent predictions. We introduce EDELINE, a novel framework that integrates diffusion models with linear-time state space modelsto enhance memory retention and temporal consistency. EDELINE employs a recurrent embedding module based on Mamba SSMs for processing unbounded sequences, a unified architecture for joint reward and termination prediction, and dynamic loss harmonization to balance multi-task learning. Our results across multiple benchmarks demonstrate EDELINE’s superiority and robustness over prior baselines in long-horizon tasks.
Machine Learning, ICML
1 Introduction
World models (Ha & Schmidhuber, 2018), which simulate environment dynamics to enable agent planning and reasoning, represent a cornerstone of modern reinforcement learning (RL). The ability to learn compressed environment representations (Hafner et al., 2024; Schrittwieser et al., 2020) enables policy optimization through imagined trajectories and improves sample efficiency (Ye et al., 2021) compared to traditional RL methods. This capability proves particularly valuable for real-world applications, such as robotics and autonomous systems. Recent advances in world modeling demonstrate remarkable success across diverse environments, including real-world tasks (Wu et al., 2022).
Existing world models can be broadly categorized into two paradigms: latent-space models and generative models. Latent-space approaches, such as (Hafner et al., 2020, 2021, 2024), leverage recurrent neural networks (RNNs) or their variants to predict future states in a compressed latent space, which enables efficient policy optimization. However, this compression often results in information loss, which compromises both generality and reconstruction quality. On the other hand, generative models, particularly diffusion-based approaches like (Alonso et al., 2024), have revolutionized world modeling by producing high-fidelity visual predictions through noise-reversal processes. Nevertheless, prior generative models rely on fixed-length observation-action windows, which truncate historical context and fail to capture extended temporal dependencies. This limitation becomes particularly problematic in partially observable environments, where agents must retain and reason over long sequences of observations to make informed decisions. Moreover, the architectural separation of reward prediction, termination signals, and observation modeling in existing frameworks can lead to suboptimal representation sharing and optimization conflicts, further hindering performance.
In order to mitigate the long sequence dependency issues, recent emergence of state space models (SSMs), such as (Gu et al., 2022a, b; Smith et al., 2023; Gu & Dao, 2024b, a) offer a complementary advantage as they are tailored to model long-term dependencies efficiently. With linear-time complexity and selective state updates (Gu & Dao, 2024b), SSMs can theoretically process unbounded sequences while preserving critical historical information. This capability is particularly valuable for world modeling, where accurate trajectory prediction often necessitates retaining and reasoning over extended observation-action histories.
Motivated by these considerations, we introduce EDELINE (Enhancing Diffusion-basEd world models via LINEar-time sequence modeling), a unified framework that synthesizes the advantages of diffusion models and SSMs. EDELINE advances the state-of-the-art through three key innovations: (1) Memory Enhancement: A recurrent embedding module (REM) based on Mamba SSMs for processing unbounded observation-action sequences to enable adaptive memory retention beyond fixed-context limitations, (2) Unified Framework: Direct conditioning of reward and termination prediction on REM hidden states for eliminating separate networks for efficient representation sharing, and (3) Dynamic Loss Harmonization: Adaptive weighting of observation, reward, and termination losses for addressing scale disparities in multi-task optimization. To validate our approach, we provide theoretical and experimental justifications to support that hidden information can preserve enhanced detail for more accurate imaginary compared with stacked frames inputs. To further demonstrate EDELINE’s capabilities, we conduct extensive validation across various benchmarks. The experimental results reveal EDELINE’s superior performance with 1.866 human normalized scores in Atari100K, which surpasses all current model-based approaches. To substantiate EDELINE’s ability to preserve long temporal information for consistent predictions, we evaluate performance on MiniGrid and VizDoom environments. Both qualitative and quantitative results demonstrate superior temporal consistency in modeling and imaginary quality across 2D and 3D environments. Our ablation studies validate the effectiveness of each architectural component for world modeling performance. Our contributions are summarized as follows:
  1. 1.
    We introduce a unified architecture that integrates a Next-Frame Predictor for future state imaginary, a Recurrent Embedding Module for temporal sequence processing, as well as a Reward/Termination Predictor.
  2. 2.
    Theoretical guarantees demonstrate that EDELINE retains strictly more information than fixed-context approaches, with advantages growing linearly over time.
  3. 3.
    Experimental results in various benchmarks validate EDELINE’s precise prediction of reward-critical elements that prior models exhibit structural inaccuracies.
2 Related Work
2.1 Diffusion Models
Diffusion models have revolutionized high-resolution image generation through their noise-reversal process. Foundational works including DDPM (Ho et al., 2020) and DDIM (Song et al., 2021a) established core principles for subsequent developments. Score-based models (Song & Ermon, 2019; Song et al., 2021b) enhanced sampling efficiency through gradient estimation of data distributions, while energy-based models (Du et al., 2023) introduced robust optimization properties through probabilistic state modeling. The application of diffusion models in reinforcement learning has expanded significantly. These models serve as policy networks for efficient offline learning (Zhendong et al., 2023; Anurag et al., 2023; Tim et al., 2023), enable diverse strategy generation in planning tasks (Michael et al., 2023; Zhixuan et al., 2023), and provide novel approaches to reward modeling (Felipe Pinto Coelho et al., 2023). MetaDiffuser (Ni et al., 2023) demonstrated their effectiveness as conditional planners in offline meta-RL, while other works have utilized diffusion models for trajectory modeling and synthetic experience generation (Cong et al., 2023).
2.2 Model-based RL and Generative Game Engines
World models serve as a fundamental component in model-based RL and enable sample-efficient and safe learning through simulated environments. SimPLe (Łukasz Kaiser et al., 2020) established the groundwork by introducing world models to the Atari domain and proposing the Atari 100k benchmark. Dreamer (Hafner et al., 2020) advanced this field through reinforcement learning from latent space predictions, which DreamerV2 (Hafner et al., 2021) further refined with discrete latents to mitigate compounding errors. DreamerV3 (Hafner et al., 2024) achieved a significant milestone by demonstrating human-level performance across multiple domains with fixed hyperparameters. Recent architectural innovations have expanded world model capabilities. TWM (Robine et al., 2023) and STORM (Zhang et al., 2023) incorporated Transformer architectures for enhanced sequence modeling, while IRIS (Micheli et al., 2023) developed a discrete image token language for structured learning. R2I (Samsami et al., 2024) addressed fundamental challenges in long-term memory and credit assignment. The integration of generative modeling approaches has further advanced world model capabilities. DIAMOND (Alonso et al., 2024) marked a significant breakthrough by incorporating diffusion models, which achieved superior visual fidelity and state-of-the-art performance on the Atari 100k benchmark. This success inspired broader applications of generative approaches in interactive environments. Another line of research explores training generative world engines using pre-collected datasets instead of RL-in-the-loop learning. GameGAN (Kim et al., 2020) pioneered this direction through GAN-based environment modeling, while Genie (Bruce et al., 2024) advanced these capabilities by generating complex platformer environments from image prompts. GameNGen (Valevski et al., 2024) established new standards for visual fidelity and scalability through diffusion-based environment simulation. It is important to note that these generative world engines do not involve RL-in-the-loop training and are therefore orthogonal to our work.
2.3 State Space Models in Reinforcement Learning
SSMs have demonstrated remarkable effectiveness in modeling long-term dependencies and structured dynamics. Foundational works (Gu et al., 2020, 2022a; Gupta et al., 2022; Gu et al., 2022b; Smith et al., 2023; Ramin et al., 2023; Gu & Dao, 2024b, a) introduced efficient sequence modeling approaches. Recent applications in RL include structured state space models for in-context learning (Chris et al., 2023), DecisionMamba (Huang et al., 2024)’s adaptation of Decision Transformer (Chen et al., 2021) architecture, and Drama (Wenlong et al., 2024)’s integration of SSMs in world model learning. Our work advances the state-of-the-art by synergistically combining diffusion-based world models with state space models. While previous works have explored these approaches separately, EDELINE demonstrates that their integration enables superior temporal consistency and extended imagination horizons in world model learning. This architectural innovation enhances the visual fidelity of diffusion models through the integration of SSMs’ efficient temporal processing capabilities, which addresses the key limitations of existing world model approaches.
3 Background
We provide additional background material on score-based diffusion models and multi-task world model learning in Appendix M. In this Section, we focus on the essential concepts necessary for understanding our EDELINE framework.
3.1 Reinforcement Learning and World Models
The problem considered in this study focuses on image-based reinforcement learning (RL), formulated as a Partially Observable Markov Decision Process (POMDP) (Åström, 1965) defined by tuple $(S,A,O,T,R,O,\gamma)$. Our formulation specifically considers high-dimensional image observations as inputs, as described in Section 1. The state space $S$ comprises states $s_{t}\in S$, while the action space $A$ can be either discrete or continuous with actions $a_{t}\in A$. The observation space $O$ contains image observations $o_{t}\in\mathbb{R}^{3\times H\times W}$. A transition function $T:S\times A\times S\rightarrow[0,1]$ characterizes the environment dynamics $p(s_{t+1}|s_{t},a_{t})$, while the reward function $R:S\times A\times S\rightarrow\mathbb{R}$ maps transitions to scalar rewards $r_{t}\in\mathbb{R}$. The observation function $O:S\times O\rightarrow[0,1]$ establishes observation probabilities $p(o_{t}|s_{t})$. The objective centers on learning a policy $\pi$ that maximizes the expected discounted return $\mathbb{E}_{\pi}[\sum_{t\geq 0}\gamma^{t}r_{t}]$, with discount factor $\gamma\in[0,1]$. Model-based Reinforcement Learning (MBRL) (Sutton, 1988) achieves this objective by learning a world model that encapsulates the environment dynamics $p(o_{t+1},r_{t}|o_{\leq t},a_{\leq t})$. MBRL enables learning in imagination through three systematic stages: (1) collecting real environment interactions, (2) updating the world model, and (3) training the policy through world model interactions.
3.2 Linear-Time Sequence Modeling with Mamba
SSMs (Gu et al., 2022a) provide an alternative paradigm to attention-based architectures for sequence modeling. The Mamba architecture (Gu & Dao, 2024b) introduces a selective state space model that offers linear time complexity and efficient parallel processing, which employs variable-dependent projection matrices to implement its selective mechanism, thus overcoming the inherent limitations of computational inefficiency and quadratic complexity in conventional SSMs (Gu et al., 2020, 2022a; Smith et al., 2023; Gu & Dao, 2024b). The foundational mechanism of Mamba is characterized by a linear continuous-time state space formulation via first-order differential equations as follows:
$\frac{\partial x(t)}{\partial t}$ $=Ax(t)+B(u(t))u(t),$ (1)
$y(t)$ $=C(u(t))x(t),$
where $x(t)$ represents the latent state, $u(t)$ denotes the input, and $y(t)$ indicates the output. The matrix $A$ adheres to specifications from (Gu et al., 2022b). The primary innovation compared to traditional SSMs lies in $B(u(t))$ and $C(u(t))$, which function as state-dependent linear operators to enable selective state updates based on input content. For discretization, the system employs the zero-order-hold (ZOH) rule (Chifu et al., 2018) to transform the $A$ and $B$ matrices into $\tilde{A}=\exp(\Delta A)$ and $\tilde{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B$, where the step size $\Delta$ serves as a variable-dependent parameter. This transformation enables SSMs to process continuous inputs as discrete signals and converts the original Linear Time-Invariant (LTI) equation into a recurrence format.
3.3 Diffusion-based World Model Learning
To adapt diffusion models for world modeling, which offers superior sample quality and tractable likelihood estimation, a key requirement is modeling the conditional distribution $p(o_{t+1}|o_{\leq t},a_{\leq t})$, where $o_{t}$ and $a_{t}$ represent observations and actions at time step $t$. The denoising process incorporates both the noised next observation and the conditioning context as input: $D_{\theta}(o_{t+1}^{\tau},\tau,o_{\leq t},a_{\leq t})$. While diffusion-based world models (Alonso et al., 2024) have shown promise, the state-of-the-art approach DIAMOND (Alonso et al., 2024) exhibits limitations although it achieves superior performance on the Atari 100k benchmark (Łukasz Kaiser et al., 2020). These models face two critical limitations. The first limitation stems from their constrained conditioning context, which typically considers only the most recent observations and actions. For instance, DIAMOND restricts its context to the last four observations and actions in the sequence. This constraint impairs the model’s capacity to capture long-term dependencies and leads to inaccurate predictions in scenarios that require extensive historical context. The second limitation in current diffusion-based world models lies in their architectural separation of predictive tasks. For example, DIAMOND implements a separate neural network with convolutional and recurrent layers for reward and termination prediction. This separation prevents the sharing of learned representations between the diffusion model and these predictive tasks and results in reduced overall learning efficiency of the system.
4 Motivational Experiments
(a) Qualitative results in MemoryS7 for DIAMOND and ours.
(b) Quantitative results in MemoryS7/S9: DIAMOND & ours.
Figure 1: Motivational examples for both qualitative and quantitative evidences to demonstrate that DIAMOND face difficulties in imagining accurate future under long horizon memorization tasks.
To substantiate the memory limitations of the DIAMOND model, we conducted experiments using the MiniGrid MemoryS7 and MemoryS9 environments (Chevalier-Boisvert et al., 2023). These experiments evaluate memory consistency and temporal prediction capabilities in world models. Fig. 1 presents qualitative comparisons among the DIAMOND model, our proposed EDELINE method, and the Oracle world model. The world models were trained on MiniGrid MemoryS7, and their state predictions were evaluated. The upper half of Fig. 1 (a) shows an initial action sequence of four consecutive left turns followed by four no-op actions. DIAMOND and EDELINE then autoregressively predicted the subsequent states. The lower half of Fig. 1 (a) reveals the prediction outcomes. At timestep 9, DIAMOND generated incorrect predictions with an extra key object in the state. The model’s predictions at timesteps 10 to 11 exhibited inaccuracies in the outer wall representations. These prediction errors result from the environment’s partial observability and DIAMOND’s four-frame context limitation. The input of four idle frames at timestep 9 led to loss of earlier state context. In contrast, EDELINE maintained accurate predictions throughout the sequence. This improved performance stems from EDELINE’s architectural design, which addresses the memory constraints inherent in DIAMOND. The accurate prediction at timestep 12 by DIAMOND can be attributed to its access to the 8th state, which enabled accurate predictions at timestep 13 through the correct 12th state context. Fig. 1 (b) illustrates DIAMOND’s performance limitations in both MemoryS7 and MemoryS9 environments due to insufficient memory capacity in partially observable scenarios. In contrast, EDELINE shows near-optimal performance under these conditions.
5 Methodology
Figure 2: Framework Overview of EDELINE.
The model integrates three principal components: (1) An U-Net-like Next-Frame Predictor enhanced by adaptive group normalization and cross-attention mechanisms, (2) A Recurrent Embedding Module built on Mamba architecture for temporal sequence processing, and (3) A Reward/Termination Predictor implemented through linear layers. The EDELINE framework uses shared hidden representations across the components for efficient world model learning.
Conventional diffusion-based world models (Alonso et al., 2024) demonstrate promise in learning environment dynamics yet face fundamental limitations in memory capacity and horizon prediction consistency. To address these challenges, this paper presents EDELINE, as illustrated in Fig 2, a unified architecture that integrates state space models (SSMs) with diffusion-based world models. EDELINE’s core innovation lies in its integration of SSMs for encoding sequential observations and actions into hidden embeddings, which a diffusion model then processes for future frame prediction. This hybrid design maintains temporal consistency while generating high-quality visual predictions. A Convolutional Neural Network based actor processes these predicted frames to determine actions, thus enabling autoregressive generation of imagined trajectories for policy optimization.
5.1 World Model Learning
The core architecture of EDELINE consists of a Recurrent Embedding Module (REM) $f_{\phi}$ that processes the history of observations and actions $(o_{0},a_{0},o_{1},a_{1},...,o_{t},a_{t})$ to generate a hidden embedding $h_{t}$ through recursive computation. This embedding enables the Next-Frame Predictor $p_{\phi}$ to generate predictions of the subsequent observation $\hat{o}_{t+1}$. The architecture further incorporates dedicated Reward and Termination Predictors to estimate the reward $\hat{r}_{t}$ and episode termination signal $\hat{d}_{t}$ respectively. The trainable components of EDELINE’s world model are formalized as:
  • Recurrent Embedding Module: $h_{t}=f_{\phi}(h_{t-1},o_{t},a_{t})$
  • Next-Frame Predictor: $\hat{o}_{t+1}\sim p_{\phi}(\hat{o}_{t+1}|h_{t})$
  • Reward Predictor: $\hat{r}_{t}\sim p_{\phi}(\hat{r}_{t}|h_{t})$
  • Termination Predictor: $\hat{d}_{t}\sim p_{\phi}(\hat{d}_{t}|h_{t})$
5.1.1 Recurrent Embedding Module
While DIAMOND, the current state-of-the-art in diffusion-based world models, relies on a fixed context window of four previous observations and actions sequence, the proposed EDELINE architecture advances beyond this limitation through a recurrent architecture for extended temporal sequence processing. Specifically, we provide theoretical evidence in Theorem F.1 to support that recurrent embedding can preserve more information compared with stacked frames inputs. At each timestep $t$, the Recurrent Embedding Module processes the current observation-action pair $(o_{t},a_{t})$ to update a hidden state $h_{t}=f_{\phi}(h_{t-1},o_{t},a_{t})$. The implementation of REM utilizes Mamba (Gu & Dao, 2024b), an SSM architecture that offers distinct advantages for world modeling. This architectural selection is motivated by the limitations of current sequence processing methods in deep learning. Self-attention-based Transformer architectures, despite their strong modeling capabilities, suffer from quadratic computational complexity which impairs efficiency. Traditional recurrent architectures including Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Chung et al., 2014) experience gradient instability issues that affect dependency learning. In contrast, SSMs provide an effective alternative through linear-time sequence processing coupled with robust memorization capabilities via their state-space formulation. The adoption of Mamba emerges as a promising choice due to its demonstrated effectiveness in modeling temporal patterns across various sequence modeling tasks. Section 6.3 presents a comprehensive ablation study that evaluates different architectural choices for the REM.
5.1.2 Next-Frame Predictor
While motivated by DIAMOND’s success in diffusion-based world modeling, EDELINE introduces significant architectural innovations in its Next-Frame Predictor to enhance temporal consistency and feature integration. At time step $t$, the model conditions on both the last $L$ frames and the hidden embedding $h_{t}$ from the Recurrent Embedding Module to predict the next frame $\hat{o}_{t+1}$. The predictive distribution $p_{\phi}(o^{0}_{t+1}|h_{t})$ is implemented through a denoising diffusion process, where $D_{\phi}$ functions as the denoising network. Let $y_{t}^{\tau}=(\tau,o^{0}_{t-L+1},...,o^{0}_{t},h_{t})$ represent the conditioning information, where $\tau$ represents the diffusion time. The denoising process can be formulated as $o^{0}_{t+1}=D_{\phi}(o^{\tau}_{t+1},y_{t}^{\tau}).$ To effectively integrate both visual and hidden information, $D_{\phi}$ employs two complementary conditioning mechanisms. First, the architecture incorporates (Zheng et al., 2020) layers within each residual block to condition normalization parameters on the hidden embedding $h_{t}$ and diffusion time $\tau$, which establishes context-aware feature normalization (Zheng et al., 2020). This design significantly extends DIAMOND’s implementation, which limits AGN conditioning to $\tau$ and action embeddings only. The second key innovation introduces cross-attention blocks inspired by Latent Diffusion Models (LDMs), which utilize $h_{t}$ and $\tau$ as context vectors. The UNet’s feature maps generate the query, while $h_{t}$ and $\tau$ project to keys and values. This novel attention mechanism, which is absent in DIAMOND, facilitates the fusion of spatial-temporal features with abstract dynamics encoded in $h_{t}$. The observation modeling loss $\mathcal{L}_{\text{obs}}(\phi)$ is defined based on Eq. (45), and can be formulated as follows:
$\mathcal{L}_{\text{obs}}(\phi)=\mathbb{E}\left[\|D_{\phi}(o^{\tau}_{t+1},y_{t}^{\tau})-o^{0}_{t+1}\|^{2}\right].$ (2)
5.1.3 Reward / Termination Predictor
EDELINE advances beyond DIAMOND’s architectural limitations through an integrated approach to reward and termination prediction. Rather than employing separate neural networks, EDELINE leverages the rich representations from its REM. The reward and termination predictors are implemented as multilayer perceptrons (MLPs) that utilize the deterministic hidden embedding $h_{t}$ as their conditioning input. This architectural unification enables efficient representation sharing across all predictive tasks. EDELINE processes both reward and termination signals as probability distributions conditioned on the hidden embedding: $p_{\phi}(\hat{r}_{t}|h_{t})$ and $p_{\phi}(\hat{d}_{t}|h_{t})$ respectively. The predictors are optimized via negative log-likelihood losses, expressed as:
$\mathcal{L}_{\text{rew}}(\phi)=-\ln p_{\phi}(r_{t}|h_{t}),\mathcal{L}_{\text{end}}(\phi)=-\ln p_{\phi}(d_{t}|h_{t}).$ (3)
This unified architectural design represents an improvement over DIAMOND’s separate network approach, where reward and termination predictions require independent representation learning from the world model. The integration of these predictive tasks with shared representations enables REM to learn dynamics that encompass all relevant aspects of the environment. The architectural efficiency facilitates enhanced learning effectiveness and better performance.
Table 1: Game scores and overall human-normalized scores on the $26$ games in the Atari $100$k benchmark. Results are averaged over 3 seeds, with bold numbers indicating the best performing method for each metric.
Game Random Human SimPLe TWM IRIS STORM DreamerV3 Drama DIAMOND EDELINE (ours)
Alien 227.8 7127.7 616.9 674.6 420.0 983.6 959.4 820 744.1 974.6
Amidar 5.8 1719.5 74.3 121.8 143.0 204.8 139.1 131 225.8 299.5
Assault 222.4 742.0 527.2 682.6 1524.4 801.0 705.6 539 1526.4 1225.8
Asterix 210.0 8503.3 1128.3 1116.6 853.6 1028.0 932.5 1632 3698.5 4224.5
BankHeist 14.2 753.1 34.2 466.7 53.1 641.2 648.7 137 19.7 854.0
BattleZone 2360.0 37187.5 4031.2 5068.0 13074.0 13540.0 12250.0 10860 4702.0 5683.3
Boxing 0.1 12.1 7.8 77.5 70.1 79.7 78.0 78 86.9 88.1
Breakout 1.7 30.5 16.4 20.0 83.7 15.9 31.1 7 132.5 250.5
ChopperCommand 811.0 7387.8 979.4 1697.4 1565.0 1888.0 410.0 1642 1369.8 2047.3
CrazyClimber 10780.5 35829.4 62583.6 71820.4 59324.2 66776.0 97190.0 52242 99167.8 101781.0
DemonAttack 152.1 1971.0 208.1 350.2 2034.4 164.6 303.3 201 288.1 1016.1
Freeway 0.0 29.6 16.7 24.3 31.1 33.5 0.0 15 33.3 33.8
Frostbite 65.2 4334.7 236.9 1475.6 259.1 1316.0 909.4 785 274.1 286.8
Gopher 257.6 2412.5 596.8 1674.8 2236.1 8239.6 3730.0 2757 5897.9 6102.3
Hero 1027.0 30826.4 2656.6 7254.0 7037.4 11044.3 11160.5 7946 5621.8 12780.8
Jamesbond 29.0 302.8 100.5 362.4 462.7 509.0 444.6 372 427.4 784.3
Kangaroo 52.0 3035.0 51.2 1240.0 838.2 4208.0 4098.3 1384 5382.2 5270.0
Krull 1598.0 2665.5 2204.8 6349.2 6616.4 8412.6 7781.5 9693 8610.1 9748.8
KungFuMaster 258.5 22736.3 14862.5 24554.6 21759.8 26182.0 21420.0 17236 18713.6 31448.0
MsPacman 307.3 6951.6 1480.0 1588.4 999.1 2673.5 1326.9 2270 1958.2 1849.3
Pong -20.7 14.6 12.8 18.8 14.6 11.3 18.4 15 20.4 20.5
PrivateEye 24.9 69571.3 35.0 86.6 100.0 7781.0 881.6 90 114.3 99.5
Qbert 163.9 13455.0 1288.8 3330.8 745.7 4522.5 3405.1 796 4499.3 6776.2
RoadRunner 11.5 7845.0 5640.6 9109.0 9614.6 17564.0 15565.0 14020 20673.2 32020.0
Seaquest 68.4 42054.7 683.3 774.4 661.3 525.2 618.0 497 551.2 2140.1
UpNDown 533.4 11693.2 3350.3 15981.7 3546.2 7985.0 7567.1 7387 3856.3 5650.3
#Superhuman (↑) 0 N/A 1 8 10 10 9 7 11 13
Mean (↑) 0.000 1.000 0.332 0.956 1.046 1.266 1.124 0.989 1.459 1.866
Median (↑) 0.000 1.000 0.134 0.505 0.289 0.580 0.485 0.270 0.373 0.817
IQM (↑) 0.000 1.000 0.130 0.459 0.501 0.636 0.487 - 0.641 0.940
Optimality Gap (↓) 1.000 0.000 0.729 0.513 0.512 0.433 0.510 - 0.480 0.387
5.1.4 EDELINE World Model Training
The world model integrates an innovative end-to-end training strategy with a self-supervised approach. EDELINE extends the harmonization technique from HarmonyDream (Ma et al., 2024) through the adoption of harmonizers $w_{o}$ and $w_{r}$, which dynamically balance the observation modeling loss $\mathcal{L}_{\text{obs}}(\phi)$ and reward modeling loss $\mathcal{L}_{\text{rew}}(\phi)$. This adaptive mechanism results in the total loss function $\mathcal{L}(\phi)$:
$\begin{split}\mathcal{L}(\phi)=w_{0}\mathcal{L}_{\text{obs}}(\phi)&+w_{r}\mathcal{L}_{\text{rew}}(\phi)+\mathcal{L}_{\text{end}}(\phi)\\ &+\log(w_{o}^{-1})+\log(w_{r}^{-1})\end{split}$ (4)
To optimize computational efficiency while ensuring robust learning, the Next-Frame Predictor learns to utilize hidden embeddings from any timestep through strategic random sampling for $\mathcal{L}_{\text{obs}}(\phi)$. For a sequence of length $T$, it follows:
$\mathcal{L}_{\text{obs}}(\phi)=\|\hat{o}^{0}_{t+1}-o_{t+1}\|^{2},$ (5)
where $i\sim\text{Uniform}\{1,2,\ldots,T-1\},\quad\hat{o}^{0}_{t+1}\sim p(\hat{o}^{0}_{t+1}\mid h_{t}).$ For reward and termination prediction, EDELINE utilizes cross-entropy losses averaged over the sequence, which can be formulated as: $\mathcal{L}_{\text{rew}}(\phi)=\frac{1}{T}\sum_{t=1}^{T}\text{CrossEnt}(\hat{r}_{t},r_{t}),$ $\mathcal{L}_{\text{end}}(\phi)=\frac{1}{T}\sum_{t=1}^{T}\text{CrossEnt}(\hat{d}_{t},d_{t}).$ This unified training approach, combining random sampling strategies with dynamic loss harmonization, demonstrates superior efficiency compared to DIAMOND’s separate network methodology, as validated in our results presented in Section 6. Moreover, the quantitative analysis presented in Appendix G reveals substantial reductions in world model training duration.
5.2 Agent Behavior Learning
To enable fair comparison and demonstrate the effectiveness of EDELINE’s world model architecture, the agent architecture adopts the same optimization framework as DIAMOND. Specifically, the agent integrates policy $\pi_{\theta}$ and value $V_{\theta}$ networks with REINFORCE value baseline and Bellman error optimization using $\lambda$-returns (Alonso et al., 2024). The training framework executes a procedure with three key phases: experience collection, world model updates, and policy optimization. This method, as formalized in Algorithm 1, follows the established paradigms in model-based RL literature (Łukasz Kaiser et al., 2020; Hafner et al., 2020; Micheli et al., 2023; Alonso et al., 2024). To ensure reproducibility, we provide extensive details in the Appendix, with documentation of objective functions, and the hyperparameter configurations in Appendices J, I, respectively.
6 Experiments
This section presents our experimental results of EDELINE. We provide details of our experimental setup in Appendix E.
6.1 Atari 100k Experiments
Following standard evaluation paradigms for world models, we evaluate EDELINE on the Atari 100k benchmark. For performance quantification, we adopt the human-normalized score (HNS) (Wang et al., 2016), which measures agent performance relative to human and random baselines:
normed score=agent scorerandom scorehuman scorerandom $\text{normed score}=\frac{\text{agent score}-\text{random score}}{\text{human score}-\text{random score}}$ (6)
Figure 3: Mean, median and interquartile HNS.
Fig. 3 presents stratified bootstrap confidence intervals following Agarwal et al. (2021)’s recommendations for point estimate limitations. It can be observed that EDELINE exhibits exceptional performance across this benchmark. Our approach surpasses human players in 13 games with a superhuman mean HNS of 1.87, median of 0.82, and IQM of 0.94. These metrics demonstrate superior performance compared to existing model-based reinforcement learning baselines without look-ahead search techniques. For detailed quantitative analysis, Table 1 provides comprehensive scores for all games. The superior performance of EDELINE stems from its ability to preserve the visual fidelity advantages of diffusion-based world models. It architecture demonstrates significant improvements over DIAMOND in memory-intensive environments such as BankHeist and Hero. The enhanced performance arises from EDELINE’s memorization capabilities, combined with harmonizer-enabled precise visual detail selection for reward-relevant features. Appendices A, B, and C provide additional training curves, qualitative analyses, and performance metrics, respectively.
6.2 ViZDoom Experiments
To evaluate EDELINE’s memory retention and imagination consistency in more challenging scenarios, we conduct experiments in the first-person 3D environment ViZDoom, which presents increased complexity compared to Atari 100k’s two-dimensional perspective. Our evaluation encompasses five default scenarios (Basic, DeadlyCorridor, DefendCenter, HealthGathering, PredictPosition), with detailed scenario descriptions and reward configurations available in Appendix L. The experimental results demonstrate EDELINE’s superior performance over DIAMOND in scenarios that demand sophisticated 3D scene modeling (DeadlyCorridor) and spatiotemporal prediction (PredictPosition). To qualitatively validate the enhanced memorization and imagination capabilities of EDELINE in 3D environments, we provide qualitative visualization in Figs. 4 and 5, which illustrate EDELINE’s accurate environment detail inference in cases where DIAMOND exhibits prediction failures.
Figure 4: Qualitative comparison on the DeadlyCorridor scenario. Each row shows ground truth, EDELINE’s predictions, and DIAMOND’s predictions, respectively. Colored boxes highlight successful (green) and failed (red) predictions of task-critical visual elements including enemies, particle effects from hits, and the armor. EDELINE accurately predicts the reward-relevant visual details. DIAMOND captures only basic environment structure.
Figure 5: Qualitative comparison on PredictPosition scenario. Each row shows ground truth, EDELINE’s predictions, and DIAMOND’s predictions. Numbers in boxes indicate predicted rewards, where 1 denotes a successful hit on the enemy. EDELINE correctly anticipates the reward while preserving visual fidelity.
6.3 Ablation Studies
6.3.1 Environments
The ablation studies in Sections 6.3.2 and 6.3.3 focus on five representative environments for validating the proposed key components. These include four Atari games where EDELINE demonstrates significant improvements over DIAMOND (BankHeist, DemonAttack, Hero, Seaquest), and MiniGrid-MemoryS9 for memory capability evaluation. This selection provides comprehensive validation across visual prediction quality and memorization requirements.
6.3.2 Choice of REM architecture
To validate the selection of Mamba for REM, we compare its performance against traditional linear-time sequence models GRU and LSTM across five environments. Fig. 6 illustrates that while all models achieve reasonable performance, Mamba demonstrates more stable learning curves and superior final performance, especially in memory-intensive tasks such as BankHeist and MiniGrid-MemoryS9. GRU and LSTM models exhibit increased training variance with lower final scores, which validates Mamba’s effectiveness.
Figure 6: Performance comparison of different linear-time sequence models as REM architecture across five environments. Training curves show mean and standard deviation over three seeds. Mamba (red) shows more stable training progression and superior final performance compared to GRU (blue) and LSTM (green).
Figure 7: Ablation study comparing EDELINE with cross-attention blocks (blue) and without (green) across five test environments. Training curves depict mean and standard deviation over three seeds. EDELINE’s cross-attention mechanism provides advantages in environments requiring rich contextual information processing.
Figure 8: Ablation study of harmonizer effectiveness on the Atari 100k benchmark. The bars depict mean HNS for DreamerV3 baseline (1.124), HarmonyDream (1.380, +22.8% over DreamerV3), EDELINE without harmonizers (1.674), and full EDELINE (1.866, +11.5% improvement from harmonizers). Detailed per-game scores and qualitative analysis are provided in Appendices D
6.3.3 Cross-attention in Next-Frame Predictor
To evaluate whether cross-attention improves the Next-Frame Predictor’s ability to process information-rich hidden embeddings, we examine EDELINE with and without this mechanism. As depicted in Fig. 7, the original EDELINE shows superior performance in BankHeist, MemoryS9, and Seaquest where rich contextual information processing proves essential. The MemoryS9 environment validates this necessity, as models must integrate historical information for complete representation reconstruction. Cross-attention enables effective fusion of hidden embeddings with visual features for temporal context integration.
6.3.4 Effect of Harmonizers
To validate the effectiveness of harmonizers, we evaluate performance across the full Atari 100k benchmark. The experimental results in Fig. 8 illustrate that harmonizer integration enhances DreamerV3 (HarmonyDream) performance from 1.124 to 1.380 mean HNS. Our EDELINE architecture achieves further substantial gains with harmonizer incorporation, advancing from 1.674 to 1.866 mean HNS (+11.5%). The analysis shows that while EDELINE establishes strong performance without harmonizers (1.674 mean HNS), harmonizer integration elevates performance to 1.866 mean HNS. This improvement validates the effectiveness of dynamic balance between visual and reward model learning.
7 Conclusions
In this work, we addressed the limitations of current diffusion-based world models in handling long-term dependencies and maintaining prediction consistency. Through the integration of Mamba SSMs, EDELINE effectively processed extended observation-action sequences through its recurrent embedding module, which enabled adaptive memory retention beyond fixed-context approaches. The unified framework eliminated architectural separation between observation, reward, and termination prediction, which fostered efficient representation sharing. Dynamic loss harmonization further mitigated optimization conflicts arising from multi-task learning. Extensive experiments on Atari100K, MiniGrid, and VizDoom validated EDELINE’s state-of-the-art performance, with significant improvements in quantitative metrics and qualitative prediction fidelity.
Appendix A Atari 100K training curves
Figure 9: Training curves of EDELINE on the Atari100k benchmark for individual games (400K environment steps). The solid lines represent the average scores over 3 seeds, and the filled areas indicate the standard deviation across these 3 seeds.
Appendix B Atari 100k Qualitative Analysis
To provide deeper insights into EDELINE’s superior performance, we conduct qualitative analysis on three Atari games where our approach demonstrates the most significant improvements over DIAMOND: BankHeist, DemonAttack, and Hero. Figure 10 presents temporal sequences comparing ground truth gameplay with predictions from both EDELINE and DIAMOND world models.
In BankHeist, agents maximize scores through repeated map traversal to encounter new enemies. EDELINE’s SSM-enhanced world model maintains consistent tracking of the player character position throughout prediction sequences, while DIAMOND’s model shows progressive degradation with the character eventually disappearing from predictions. For DemonAttack, EDELINE successfully captures the relationship between enemy hits and score updates in its predictions. DIAMOND preserves basic visual structure but fails to reflect these crucial state transitions. The Hero environment showcases EDELINE’s long-range prediction capabilities, accurately capturing sequences of the player breaking obstacles and navigating new areas. These qualitative results highlight how EDELINE’s architectural innovations - SSM-based memory, unified training, and diffusion modeling - enable robust state tracking, action-consequence modeling, and temporal consistency. These capabilities directly contribute to EDELINE’s superior performance on the Atari 100k benchmark.
Figure 10: Qualitative comparison of world model predictions on three Atari games. Each panel shows temporal sequences comparing ground truth with EDELINE and DIAMOND predictions. In BankHeist (top), EDELINE successfully tracks the player character while DIAMOND loses this information. DemonAttack (middle) demonstrates EDELINE’s accurate modeling of score updates upon successful hits. Hero (bottom) showcases EDELINE’s ability to maintain consistent predictions of complex character-environment interactions across extended sequences. Colored boxes highlight successful (green) and failed (red) predictions of key game elements.
Appendix C Extended Performance Analysis
(a) Performance profiles
(b) Optimality gaps
Figure 11: Additional performance analyses. (a) Performance profiles showing fraction of runs achieving scores above different human normalized score thresholds. (b) Optimality gaps demonstrating relative distance to theoretical optimal performance across different methods.
To provide additional performance insights, we analyze EDELINE using performance profiles and optimality gaps (Agarwal et al., 2021). Fig. 11(a) shows the empirical cumulative distribution of human-normalized scores (HNS) across all 26 Atari games. The y-axis indicates the fraction of games achieving scores above each HNS threshold $\tau$. EDELINE consistently maintains a higher fraction of games across different thresholds compared to baseline methods, particularly in the middle range ($\tau$ between 1 and 4). We further analyze model performance through optimality gaps shown in Fig. 11(b). The optimality gap measures the distance between model performance and theoretical optimal behavior. EDELINE achieves the smallest optimality gap among all compared methods, demonstrating its effectiveness in approaching optimal performance across the Atari 100k benchmark suite. These detailed analyses complement the aggregate metrics presented in Section 6.1 by providing a more granular view of performance distribution and theoretical efficiency.
Figure 12: Comparison of state-of-the-art model-based RL methods without using look-ahead search techniques on the Atari 100k benchmark. EDELINE outperforms all existing model-based approaches on the Atari 100k benchmark. Previous methods can be categorized by their world model architectures: Transformer-based models (TWM (Robine et al., 2023), IRIS (Micheli et al., 2023), REM (Cohen et al., 2024), STORM (Zhang et al., 2023), $\Delta$-IRIS (Alonso et al., 2023), TWISTER (Anonymous, 2025)), RNN-based models (DreamerV3 (Hafner et al., 2024), PaMoRL (Wang et al., 2024), HarmonyDream (Ma et al., 2024)), SSM-based models (Drama (Wenlong et al., 2024)), and Diffusion-based models (DIAMOND (Alonso et al., 2024)). Our proposed EDELINE advances the state-of-the-art by integrating diffusion modeling with state space models, combining their respective strengths in visual generation and temporal modeling.
Appendix D Harmonizers Ablation Results
(a) BankHeist
(b) RoadRunner
Figure 13: Effect of harmonizers on world model predictions. (a) Without harmonizers, EDELINE fails to maintain consistent character modeling in BankHeist. (b) In RoadRunner, removal of harmonizers leads to loss of reward-relevant visual details. Colored boxes highlight successful (green) and failed (red) predictions of key game elements.
Table 2: To evaluate the benefits of Harmonizers in detail, we compare DreamerV3 and its variant with Harmonizers (HarmonyDream), alongside EDELINE without Harmonizers and the complete EDELINE on the $26$ games in the Atari $100$k benchmark. Bold numbers indicate the highest scores.
Game Random Human DreamerV3 HarmonyDream EDELINE w/o Harmonizers EDELINE (ours)
Alien 227.8 7127.7 959.4 889.7 1086.2 974.6
Amidar 5.8 1719.5 139.1 141.1 212.5 299.5
Assault 222.4 742.0 705.6 1002.7 1180.1 1225.8
Asterix 210.0 8503.3 932.5 1140.3 4612.8 4224.5
BankHeist 14.2 753.1 648.7 1068.6 139.8 854.0
BattleZone 2360.0 37187.5 12250.0 16456.0 2685.0 5683.3
Boxing 0.1 12.1 78.0 79.6 88.3 88.1
Breakout 1.7 30.5 31.1 52.6 255.7 250.5
ChopperCommand 811.0 7387.8 410.0 1509.6 2616.5 2047.3
CrazyClimber 10780.5 35829.4 97190.0 82739.0 96889.5 101781.0
DemonAttack 152.1 1971.0 303.3 202.6 924.2 1016.1
Freeway 0.0 29.6 0.0 0.0 33.8 33.8
Frostbite 65.2 4334.7 909.4 678.7 289.0 286.8
Gopher 257.6 2412.5 3730.0 13042.8 7982.7 6102.3
Hero 1027.0 30826.4 11160.5 13378.0 9366.2 12780.8
Jamesbond 29.0 302.8 444.6 317.1 527.5 784.3
Kangaroo 52.0 3035.0 4098.3 5117.6 3970.0 5270.0
Krull 1598.0 2665.5 7781.5 7753.6 8762.7 9748.8
KungFuMaster 258.5 22736.3 21420.0 22274.0 14088.5 31448.0
MsPacman 307.3 6951.6 1326.9 1680.7 1773.3 1849.3
Pong -20.7 14.6 18.4 18.6 18.4 20.5
PrivateEye 24.9 69571.3 881.6 2932.2 73.0 99.5
Qbert 163.9 13455.0 3405.1 3932.5 5062.3 6776.2
RoadRunner 11.5 7845.0 15565.0 14646.4 23272.5 32020.0
Seaquest 68.4 42054.7 618.0 665.3 1277.2 2140.1
UpNDown 533.4 11693.2 7567.1 10873.6 2844.6 5650.3
#Superhuman (↑) 0 N/A 9 11 11 13
Mean (↑) 0.000 1.000 1.124 1.364 1.674 1.866
Table 2 presents quantitative performance comparison between EDELINE with and without harmonizers. The ablation results demonstrate significant performance degradation across multiple environments when harmonizers are removed, with particularly notable drops in games requiring precise reward related visual detail retention (BankHeist and RoadRunner). Qualitative analysis on these two environments, which showed the largest performance improvements with harmonizers, reveals how harmonizers contribute to world model performance. In BankHeist (Fig. 13(a)), removing harmonizers causes the world model to lose track of game characters, failing to maintain consistent agent representation across prediction sequences. Similarly, in RoadRunner (Fig. 13(b)), the model without harmonizers fails to capture reward-relevant visual details, degrading its ability to model critical game state information.
These results demonstrate how harmonizers help achieve a crucial balance between observation modeling and reward prediction. Without harmonizers, the world model struggles with fine-grained task-relevant observations in both environments - tracking small character sprites in BankHeist and capturing critical game state details in RoadRunner. By maintaining this dynamic equilibrium between observation and reward modeling, harmonizers enable EDELINE to effectively learn compact task-centric dynamics while preserving essential visual details for sample-efficient learning.
Appendix E Experimental Setup
We evaluate EDELINE on the Atari 100k benchmark (Schwarzer et al., 2023), which serves as the standard evaluation protocol in recent model-based RL literature for fair comparison. In addition, our experimental validation extends to ViZDoom (Kempka et al., 2016) and MiniGrid (Chevalier-Boisvert et al., 2023) environments to demonstrate broader applicability. To ensure statistical significance, all reported results represent averages across three independent runs. The Atari 100k benchmark (Schwarzer et al., 2023) encompasses 26 diverse Atari games that evaluate various aspects of agent capabilities. Each agent receives a strict limitation of 100k environment interactions for learning, in contrast to conventional Atari agents that typically require 50 million steps. EDELINE’s performance is evaluated against current state-of-the-art world model-based approaches, including DIAMOND (Alonso et al., 2024), STORM (Zhang et al., 2023), DreamerV3 (Hafner et al., 2024), IRIS (Micheli et al., 2023), TWM (Robine et al., 2023), and Drama (Wenlong et al., 2024). For evaluating 3D scene understanding capabilities, we employ VizDoom scenarios that demand sophisticated 3D spatial reasoning in first-person environments. This provides a crucial testing ground beyond the third-person perspective of Atari environments. Furthermore, the MiniGrid memory scenarios evaluate memorization capabilities through tasks that require information retention across extended time horizons.
Appendix F Theoretical Analysis of Mamba SSM Information Retention
F.1 Background and Motivation
Traditional deep reinforcement learning approaches rely on stacked frame inputs $s_{t}=[o_{t-k},\ldots,o_{t}]$ to capture temporal dependencies. This approach presents two fundamental limitations:
  1. 1.
    Fixed Memory Window: Information horizon is strictly bounded by window size $k$
  2. 2.
    Uniform Treatment: All observations within the window receive equal weighting regardless of importance
We propose replacing this architecture with Mamba State Space Models (SSM) utilizing hidden states $h_{t}$. Our analysis demonstrates superior information retention properties through rigorous mathematical proof.
F.2 Main Theorem and Implications
Theorem F.1 (Information Retention Superiority).
For any observation sequence $H_{t}=\{o_{1},\ldots,o_{t}\}$, with stacked frames $s_{t}$ of window size $k$ and Mamba SSM hidden state $h_{t}$, the following inequality holds:
$I(h_{t};H_{t})\geq I(s_{t};H_{t})+\Delta(t,k)$ (7)
where $\Delta(t,k)=t\epsilon-(k+1)H(o)$ becomes strictly positive when $t>\frac{(k+1)H(o)}{\epsilon}$.
This theorem establishes that Mamba’s information retention capability strictly surpasses traditional stacked frame approaches after a critical threshold, with the advantage growing linearly with sequence length. The term $\epsilon$ represents the minimum information retention rate per timestep, while $H(o)$ denotes the entropy of a single observation. The critical threshold $\frac{(k+1)H(o)}{\epsilon}$ represents the sequence length at which Mamba’s advantages begin to manifest.
F.3 Detailed Proof
The proof proceeds in three parts:
  1. 1.
    First, we establish an upper bound on information content in stacked frames
  2. 2.
    Then, we prove a lower bound on Mamba’s information retention
  3. 3.
    Finally, we combine these results to demonstrate the superiority condition
F.3.1 Part 1: Analysis of Stacked Frame Information Content
We begin by analyzing the information content of traditional stacked frames:
Definition F.2 (Stacked Frame Representation).
The stacked frame representation $s_{t}$ is defined as:
$s_{t}=[o_{t-k},\ldots,o_{t}]\in\mathcal{O}^{k+1}$ (8)
where $\mathcal{O}$ is the observation space.
Lemma F.3 (Stacked Frame Information Bound).
The mutual information between stacked frames and history is bounded by:
$I(s_{t};H_{t})\leq(k+1)H(o)$ (9)
Proof.
We apply the chain rule of entropy and use the properties of mutual information:
$I(s_{t};H_{t})$ $\leq H(s_{t})$ (Information bound) (10)
$=H(o_{t-k},\ldots,o_{t})$ (Definition of $s_{t}$) (11)
$\leq\sum_{i=0}^{k}H(o_{t-i})$ (Chain rule) (12)
$=(k+1)H(o)$ (Stationarity) (13)
The equality in the last step holds even the observations are dependent. This demonstrates the fundamental limitation of fixed-window approaches, as the information content is strictly bounded by the window size. ∎
Proposition F.4 (Information Loss Ratio).
The stacked frame architecture exhibits an information loss ratio of:
$L_{\text{stack}}=1-\frac{I(s_{t};H_{t})}{H(H_{t})}\geq 1-\frac{k+1}{t}$ (14)
Proof.
The loss ratio can be derived as follows:
$L_{\text{stack}}$ $=1-\frac{I(s_{t};H_{t})}{H(H_{t})}$ (Definition) (15)
$\geq 1-\frac{(k+1)H(o)}{tH(o)}$ (Using Lemma 1.3) (16)
$=1-\frac{k+1}{t}$ (Simplification) (17)
This reveals that the information loss grows linearly with the sequence length $t$, demonstrating the inefficiency of fixed-window approaches for long sequences. ∎
F.3.2 Part 2: Mamba SSM Information Retention Analysis
Definition F.5 (Mamba SSM State Update).
The Mamba SSM implements recursive state updates through:
$h_{t}=A_{t}h_{t-1}+B_{t}o_{t}$ (18)
where $h_{t}\in\mathbb{R}^{d}$ denotes the hidden state vector, $A_{t}\in\mathbb{R}^{d\times d}$ represents the state transition matrix, $B_{t}\in\mathbb{R}^{d\times m}$ is the input projection matrix, and $o_{t}\in\mathbb{R}^{m}$ is the observation vector.
Lemma F.6 (Information Evolution).
The mutual information between hidden state and history evolves according to:
$I(h_{t};H_{t})=I(h_{t};H_{t-1})+I(h_{t};o_{t}|H_{t-1})$ (19)
where $I(h_{t};o_{t}|H_{t-1})$ represents the new information gained from the current observation. The lemma aligns with Mamba’s recursive state update mechanism, which accumulates information adaptively.
Proof.
The proof follows from the chain rule of mutual information and the structure of the historical sequence:
$I(h_{t};H_{t})$ $=I(h_{t};\{H_{t-1},o_{t}\})$ (Definition of $H_{t}$)
$=I(h_{t};H_{t-1})+I(h_{t};o_{t}|H_{t-1})$ (Chain rule)
$=I(h_{t};H_{t-1})+\underbrace{[H(o_{t}|H_{t-1})-H(o_{t}|h_{t},H_{t-1})]}_{\text{Information gain from }o_{t}}$ (20)
The last term quantifies the reduction in uncertainty about $o_{t}$ provided by $h_{t}$, given the history $H_{t-1}$. This decomposition reveals how Mamba’s state update mechanism captures new information while preserving relevant historical context. ∎
Corollary F.7 (Information Monotonicity).
Under the Mamba update rule, the mutual information $I(h_{t};H_{t})$ is monotonically increasing in $t$:
$I(h_{t};H_{t})\geq I(h_{t-1};H_{t-1})+\epsilon_{t}$ (21)
where $\epsilon_{t}$ is the minimum information gain at time $t$.
Theorem F.8 (Information Growth).
Under Mamba’s selective memory mechanism:
$I(h_{t};H_{t})\geq\sum_{\tau=1}^{t}\epsilon_{\tau}\geq t\epsilon$ (22)
where $\epsilon=\min_{\tau}\epsilon_{\tau}>0$ is guaranteed by design.
Proof.
We proceed by induction on the sequence length $t$.
Base case $(t=1)$:
$I(h_{1};H_{1})=I(h_{1};o_{1})\geq\epsilon_{1}\geq\epsilon$ (23)
Inductive step: Assume the inductive hypothesis holds for $t-1$, i.e.,
$I(h_{t-1};H_{t-1})\geq\sum_{\tau=1}^{t-1}\epsilon_{\tau}$ (24)
Then for time step $t$:
$I(h_{t};H_{t})$ $=I(h_{t};H_{t-1})+I(h_{t};o_{t}|H_{t-1})$ (Chain rule) (25)
$\geq I(h_{t-1};H_{t-1})+\epsilon_{t}$ (DPI + Selectivity) (26)
$\geq\sum_{\tau=1}^{t-1}\epsilon_{\tau}+\epsilon_{t}=\sum_{\tau=1}^{t}\epsilon_{\tau}$ (Inductive hypothesis) (27)
$\geq t\epsilon$ (Definition of $\epsilon$) (28)
Therefore, by mathematical induction, the theorem holds for all $t\geq 1$. The linear growth in mutual information demonstrates Mamba’s ability to retain information over arbitrarily long sequences. ∎
F.3.3 Part 3: Comparative Analysis and Critical Threshold
Theorem F.9 (Information Advantage).
The information advantage $\Delta(t,k)$ of Mamba SSM over stacked frames is:
$\Delta(t,k)=I(h_{t};H_{t})-I(s_{t};H_{t})\geq t\epsilon-(k+1)H(o)$ (29)
This theorem provides practitioners with a quantitative decision tool: For any given window size $k$ and estimated retention rate $\epsilon$, one can compute the critical sequence length where Mamba becomes advantageous.
Proof.
Combining our previous results:
$\Delta(t,k)$ $=I(h_{t};H_{t})-I(s_{t};H_{t})$ (30)
$\geq t\epsilon-(k+1)H(o)$ (Theorems 2 and 3)
This advantage becomes positive when:
$t>\frac{(k+1)H(o)}{\epsilon}=t_{crit}$ (31)
F.4 Practical Implementation and Results
For practical implementation in Atari100k environments:
  • Standard window size: $k=4$
  • Empirically learned retention rate: $\epsilon\approx 0.2H(o)$
  • Resulting critical threshold: $t_{crit}=25$ frames
F.5 Optimality Conditions
The selection mechanism optimizes:
[…]
Subject to computational constraints:
$\text{rank}(A_{t})$ $\leq r$ (33)
$\|A_{t}\|_{F}$ $\leq\alpha$ (34)
where $r$ is the rank constraint and $\alpha$ bounds the Frobenius norm.
F.6 Key Theoretical Properties
  1. 1.
    Information Growth: Linear growth in mutual information with sequence length
  2. 2.
    Adaptive Memory: Input-dependent retention through learned matrices $A_{t},B_{t}$
  3. 3.
    Bounded Complexity: Linear computational scaling in sequence length
  4. 4.
    Provable Advantage: Quantifiable improvement over fixed-window approaches
These properties collectively address the core limitations of traditional approaches while maintaining computational tractability.
Appendix G Training time profile
Table 3: Detailed breakdown of training time components across different scales. Profiling performed using a Nvidia RTX 4090 with default hyperparameters. Measurements are representative, as exact durations depend on specific hardware, environment, and training stage.
Single update Time (ms) Detail (ms)
Total 548.6 148.6 + 400
   World model update 148.6 -
   Actor-Critic model update 400 15 $\times$ 24.4 + 34
    Imagination step (x 15) 24.4 17.2 + 6.5 + 0.7
     Next observation prediction 17.2 -
     Denoising step (x 3) 5.7 -
     Reward/Termination prediction 6.5 -
     Action prediction 0.7 -
    Loss computation and backward 34 -
Epoch Time (s) Detail (s)
Total 219.4 59.4 + 160
   World model 59.4 400 $\times$ 148.6 ×$\times 10^{-3}$
   Actor-Critic model 160 400 $\times$ 400 ×$\times 10^{-3}$
Run Time (days) Detail (days)
Total 2.9 2.5 + 0.4
   Training time 2.5 1000 $\times$ 219.4 / (24 $\times$ 3600)
   Other (collection, evaluation, checkpointing) 0.4 -
Table 4: Computational efficiency comparison between DIAMOND and EDELINE across different training stages, showing per-stage timing and relative differences.
Module DIAMOND (ms/s/days) EDELINE (ms/s/days) Difference (%)
Single Update
Total 543 ms 548.6 ms +1.03%
   World Model Update 203 ms 148.6 ms -26.80%
   Actor-Critic Model Update 340 ms 400 ms +17.65%
Epoch
Total 217 s 219.4 s +1.11%
   World Model 81 s 59.4 s -26.67%
   Actor-Critic Model 136 s 160 s +17.65%
Appendix H edeline algorithm
We summarize the overall training procedure of edeline in Algorithm 1 below, which is modified from Algorithm 1 in Alonso et al. (2024). We denote as $\mathcal{D}$ the replay dataset where the agent stores data collected from the real environment, and other notations are introduced in previous sections or are self-explanatory.
Algorithm 1 EDELINE
1:  function training_loop():
2:     for epochs do
3:        collect_experience(steps_collect)
4:        for steps_world_model do
5:           update_world_model()
6:        end for
7:        for steps_actor_critic do
8:           update_actor_critic()
9:        end for
10:     end for
11:  end function
12:  function collect_experience($n$):
[…]
14:     for $t=0$ to $n-1$ do
15:        Sample $a_{t}\sim\pi_{\theta}(a_{t}\mid o_{t}^{0})$
16:        $o_{t+1}^{0},r_{t},d_{t}\leftarrow\texttt{env.step}(a_{t})$
17:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{o_{t}^{0},a_{t},r_{t},d_{t}\}$
18:        if $d_{t}=1$ then
[…]
20:        end if
21:     end for
22:  end function
23:  function update_world_model():
24:     Sample indexes :=$\mathcal{I}:=\{t,\;\dots,\;t+B+H-1\}$ // burn-in $B$ steps + imagination horizon $H$
25:     Sample sequence $(o_{i}^{0},\;a_{i},\;r_{i},\;d_{i})_{i\in\mathcal{I}}\sim\mathcal{D}$
26:     Initialize $h_{t-1}$ // MAMBA hidden state
27:     for $i\in\mathcal{I}$ do
28:        $h_{i}\leftarrow f_{\phi}(o_{i}^{0},\;a_{i},\;h_{i-1})$
29:        $\hat{r}_{i}\sim p_{\phi}(\hat{r}_{i}|h_{i})$
30:        $\hat{d}_{i}\sim p_{\phi}(\hat{d}_{i}|h_{i})$
31:     end for
32:     Compute reward modeling loss: $\mathcal{L}_{\mathrm{rew}}(\phi)\;=\;\sum_{i\in\mathcal{I}}\mathrm{CE}\bigl{(}\hat{r}_{i},\;r_{i}\bigr{)}$ // CE: cross-entropy loss
33:     Compute termination signal modeling loss: $\mathcal{L}_{\mathrm{end}}(\phi)\;=\;\sum_{i\in\mathcal{I}}\mathrm{CE}\bigl{(}\hat{d}_{i},\;d_{i}\bigr{)}$ // CE: cross-entropy loss
34:     Sample index $j$ $\sim$ $\text{Uniform}\{t+B,\dots,t+B+H-1\}$ // random sampling
35:     Sample $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ // log-normal sampling from EDM
36:     Define $\tau:=\sigma$ // identity schedule from EDM
37:     Sample $o_{j}^{\tau}\sim\mathcal{N}(o_{j}^{0},\;\sigma^{2}\mathbf{I})$ // add Gaussian noise
38:     Compute $\hat{o}_{j}^{0}=D_{\phi}\bigl{(}o_{j}^{\tau},\,\tau,\,o_{j-L}^{0},\dots,o_{j-1}^{0},h_{j-1}\bigr{)}$
39:     Compute observation modeling loss $\mathcal{L}_{\mathrm{obs}}(\phi)=\|\hat{o}_{j}^{0}-o_{j}^{0}\|^{2}$
40:     Update $\phi$ according to Eq. (4)
41:  end function
42:  function update_actor_critic():
43:     Sample initial buffer $(o_{t-B+1}^{0},\;a_{t-B+1},\;\dots,\;o_{t}^{0})\sim\mathcal{D}$
44:     // Burn in LSTM states $\pi_{\theta}$, $V_{\theta}$ and MAMBA states $f_{\phi}$ with the buffer
45:     for $i=t$ to $t+H-1$ do
46:        Sample $a_{i}\sim\pi_{\theta}(a_{i}\mid o_{i}^{0})$
47:        Compute $h_{i}\leftarrow f_{\phi}(o_{i},a_{i},h_{i-1})$
48:        Sample reward $r_{i}$, next observation $o_{i+1}^{0}$, and termination $d_{i}$ via $p_{\phi}$
49:     end for
50:     Compute $V_{\theta}(o_{i}^{0})$ for $i=t,\dots,t+H$
51:     Compute RL losses $\mathcal{L}_{V}(\theta)$ and $\mathcal{L}_{\pi}(\theta)$
52:     Update $\pi_{\theta}$ and $V_{\theta}$
53:  end function
Appendix I Hyperparameters
Table 5: Hyperparameters for EDELINE.
Hyperparameter Symbol Value
General
Number of epochs 1000
Training steps per epoch 400
Environment steps per epoch 100
Batch size 32
Sequence Length $T$ 19
Epsilon (greedy) for collection $\epsilon$ 0.01
Observation shape (h,w,c) (64,64,3)
Actor Critic
Imagination horizon $H$ 15
Discount factor $\gamma$ 0.985
Entropy weight $\eta$ 0.001
$\lambda$-returns coefficient $\lambda$ 0.95
LSTM dimension 512
Residual blocks layers [1,1,1,1]
Residual blocks channels [32,32,64,64]
World Model
Number of conditioning observations $L$ 4
Burn-in length $B$ 4
Hidden embedding dimension d_model 512
Reward / Termination Model Hidden Units 512
Mamba
Mamba layers n_layers 3
Mamba state dimension d_state 16
Expand factor 2
1D Convolution Dimension d_conv 4
Residual blocks layers [1,1,1,1]
Residual blocks channels [64,64,64,64]
Action embedding dimenstion 128
Diffusion
Sampling method Euler
Number of denoising steps 3
Condition embedding dimension 256
Residual blocks layers [2,2,2,2]
Residual blocks channels [64,64,64,64]
Optimization
Optimizer AdamW
Learning rate $\alpha$ 1e-4
Epsilon 1e-8
Weight decay (World Model) 1e-2
Weight decay (Actor-Critic) 0
Max grad norm (World Model) 1.0
Max grad norm (Actor-Critic) 100.0
Appendix J Actor-Critic Learning Objectives
We follow DIAMOND (Alonso et al., 2024) in the design of our agent behavior learning. Let $o_{t}$, $r_{t}$, and $d_{t}$ denote the observations, rewards, and boolean episode terminations predicted by our world model. We denote $H$ as the imagination horizon, $V_{\theta}$ as the value network, $\pi_{\theta}$ as the policy network, and $a_{t}$ as the actions taken by the policy within the world model.
For value network training, we use $\lambda$-returns to balance bias and variance in the regression target. Given an imagined trajectory of length $H$, we define the $\lambda$-return recursively:
$\Lambda_{t}=\begin{cases}r_{t}+\gamma(1-d_{t})[(1-\lambda)V_{\theta}(o_{t+1})+\lambda\Lambda_{t+1}]&\text{if }t<H\\ V_{\theta}(o_{H})&\text{if }t=H.\end{cases}$ (35)
The value network $V_{\theta}$ is trained to minimize $\mathcal{L}_{V}(\theta)$, the expected squared difference with $\lambda$-returns over imagined trajectories:
$\mathcal{L}_{V}(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{H-1}(V_{\theta}(\mathbf{x}_{t})-\text{sg}(\Lambda_{t}))^{2}\right],$ (36)
where $\text{sg}(\cdot)$ denotes the gradient stopping operation, following standard practice (Hafner et al., 2024; Micheli et al., 2023).
For policy training, we leverage the ability to generate large amounts of on-policy trajectories in imagination using a REINFORCE objective (Sutton & Barto, 2018). The policy is trained to minimize:
$\mathcal{L}_{\pi}(\theta)=-\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{H-1}\log(\pi_{\theta}(a_{t}|o_{\leq t}))\text{sg}(\Lambda_{t}-V_{\theta}(o_{t}))+\eta\mathcal{H}(\pi_{\theta}(a_{t}|o_{\leq t}))\right],$ (37)
where $V_{\theta}(o_{t})$ serves as a baseline to reduce gradient variance, and the entropy term $\mathcal{H}$ with weight $\eta$ encourages sufficient exploration.
Appendix K EDM Network Preconditioners and Training
Following DIAMOND (Alonso et al., 2024), we use the EDM preconditioners from (Karras et al., 2022) for normalization and rescaling to improve network training:
$c_{in}^{\tau}=\frac{1}{\sqrt{\sigma(\tau)^{2}+\sigma_{data}^{2}}}$ (38)
$c_{out}^{\tau}=\frac{\sigma(\tau)\sigma_{data}}{\sqrt{\sigma(\tau)^{2}+\sigma_{data}^{2}}}$ (39)
$c_{noise}^{\tau}=\frac{1}{4}\log(\sigma(\tau))$ (40)
$c_{skip}^{\tau}=\frac{\sigma_{data}^{2}}{\sigma_{data}^{2}+\sigma^{2}(\tau)},$ (41)
where $\sigma_{data}=0.5$.
The noise parameter $\sigma(\tau)$ is sampled to maximize the effectiveness of training:
$\log(\sigma(\tau))\sim\mathcal{N}(P_{mean},P_{std}^{2}),$ (42)
where $P_{mean}=-0.4$, $P_{std}=1.2$.
Appendix L ViZDoom Environment Specifications
This appendix provides detailed specifications for the ViZDoom scenarios used in our experiments. We evaluate EDELINE on five key scenarios that test different agent capabilities in first-person 3D environments. All environments employ a frame skipping strategy of 4:1, where the agent executes each selected action for 4 consecutive frames.
L.1 DeadlyCorridor
  • Objective: Navigate through enemy fire to acquire armor at corridor end
  • Reward Mechanism:
    • +1 for each enemy killed
    • +1 for armor acquisition
    • -1 for each damage instance received
  • Evaluation Metric: Binary success if armor acquired before episode termination
  • Episode Termination: Agent death or 512 environment interaction steps.
L.2 HealthGathering
  • Objective: Survive by collecting medkits in toxic environment
  • Reward Mechanism:
    • +1 for each medkit collected
  • Evaluation Metric: Final health percentage (0-100%)
  • Episode Termination: Agent death or 512 environment interaction steps.
L.3 PredictPosition
  • Objective: Anticipate enemy movement to land delayed rocket hit
  • Reward Mechanism:
    • +1 for successful enemy kill
  • Evaluation Metric: Binary success if enemy eliminated
  • Episode Termination: Successful kill or 75 environment interaction steps.
L.4 Basic
  • Objective: Eliminate stationary enemy with limited ammunition
  • Reward Mechanism:
    • +1 for successful kill
  • Evaluation Metric: Binary success if enemy eliminated
  • Episode Termination: Successful kill or 75 environment interaction steps.
L.5 DefendCenter
  • Objective: Survive against infinite enemy waves
  • Reward Mechanism:
    • +1 per enemy killed
    • -1 for agent death
  • Evaluation Metric: Total enemies killed per episode
  • Episode Termination: Agent death or 512 environment interaction steps.
All scenarios use the following common configuration parameters unless otherwise specified:
  • Observation space: (64,64,3)
  • Action space: Discrete movement and attack actions
Appendix M Additional Background Material Section
M.1 Score-based Diffusion Generative Models
Diffusion probabilistic modeling (Sohl-Dickstein et al., 2015; Ho et al., 2020; Dhariwal & Nichol, 2021) and score-based generative modeling (Song & Ermon, 2019, 2020; Chao et al., 2022) can be unified through a forward stochastic differential equation (SDE) formulation (Song et al., 2021b). The forward diffusion process $\{{\mathbf{x}}^{\tau}\}$ with continuous time variable $\tau$ transforms the data distribution $p^{0}=p^{\text{data}}$ to prior distribution $p^{T}=p^{\text{prior}}$, expressed as:
$\mathrm{d}{\mathbf{x}}={\mathbf{f}}({\mathbf{x}},\tau)\mathrm{d}\tau+g(\tau)\mathrm{d}{\mathbf{w}},$ (43)
where ${\mathbf{f}}({\mathbf{x}},\tau)$ represents the drift coefficient, $g(\tau)$ denotes the diffusion coefficient, and ${\mathbf{w}}$ is the Wiener process. The corresponding reverse-time SDE can then be formulated as:
$\mathrm{d}{\mathbf{x}}=\left[{\mathbf{f}}({\mathbf{x}},\tau)-g(\tau)^{2}\nabla_{\mathbf{x}}\log p^{\tau}({\mathbf{x}})\right]\mathrm{d}\tau+g(\tau)\mathrm{d}\bar{{\mathbf{w}}},$ (44)
where $\bar{{\mathbf{w}}}$ is the reverse-time Wiener process. Eq. (44) enables sampling from $p^{0}$ when the (Stein) score function $\nabla_{\mathbf{x}}\log p^{\tau}({\mathbf{x}})$ is available. A common approach to estimate the score function is through the introduction of a denoiser $D_{\theta}$, which is trained to minimize the following objective:
$\mathbb{E}_{\sigma\sim p^{\text{train}}}\mathbb{E}_{{{\mathbf{x}}^{0}}\sim p^{\text{data}}}\mathbb{E}_{{\mathbf{n}}\sim\mathcal{N}(0,\sigma^{2}I)}\left[\|D_{\theta}({\mathbf{x}}^{0}+{\mathbf{n}};\sigma)-{\mathbf{x}}^{0}\|_{2}^{2}\right],$ (45)
where ${\mathbf{n}}$ is Gaussian noise with zero mean and variance determined by a variance scheduler $\sigma(\tau)$ that follows a noise distribution $p^{\text{train}}$, and $({\mathbf{x}}^{0}+{\mathbf{n}})$ corresponds to the perturbed data ${\mathbf{x}}^{\tau}$. The score function can then be estimated through: $\nabla_{{\mathbf{x}}}\log p^{\tau}({\mathbf{x}})=\frac{1}{\sigma^{2}}(D_{\theta}({\mathbf{x}};\sigma)-{\mathbf{x}})$. In practice, modeling the denoiser $D_{\theta}$ directly can be challenging due to the wide range of noise scales. To address this, EDM (Karras et al., 2022) introduces a design space that isolates key design choices, including preconditioning functions $\{c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}\}$ to modulate the unconditioned neural network $F_{\theta}$ to represent $D_{\theta}$, which can be formulated as:
$D_{\theta}({\mathbf{x}};\sigma)=c_{\text{skip}}(\sigma){\mathbf{x}}+c_{\text{out}}(\sigma)F_{\theta}(c_{\text{in}}(\sigma){\mathbf{x}};c_{\text{noise}}(\sigma)).$ (46)
The preconditioners serve distinct purposes: $c_{\text{in}}(\sigma)$ and $c_{\text{out}}(\sigma)$ maintain unit variance for network inputs and outputs across noise levels, $c_{\text{noise}}(\sigma)$ provides transformed noise level conditioning, and $c_{\text{skip}}(\sigma)$ adaptively balances signal mixing. This principled framework improves the robustness and efficiency of diffusion models, enabling state-of-the-art performance across various generative tasks.
M.2 Multi-task Essence of World Model Learning
Modern world models (Łukasz Kaiser et al., 2020; Hafner et al., 2021, 2024; Alonso et al., 2024) typically address two fundamental prediction tasks: the modeling of environment dynamics through observations and the prediction of reward signals. The learning of these tasks requires distinct considerations based on the complexity of the environment. In simple low-dimensional settings, separate learning approaches suffice for each task. However, the introduction of high-dimensional visual inputs fundamentally alters this paradigm, as partial observability creates an inherent coupling between state estimation and reward prediction. This coupling necessitates joint learning through shared representations, an approach that aligns with established multi-task learning principles (Caruana, 1997). The implementation of such joint learning through shared representations introduces several technical challenges. The integration of multiple learning objectives requires careful consideration of their relative importance and interactions. A fundamental difficulty stems from the inherent scale disparity between high-dimensional visual observations and scalar reward signals. This disparity manifests in the world model learning objective, which combines observation modeling $\mathcal{L}_{o}(\theta)$, reward modeling $\mathcal{L}_{r}(\theta)$, and dynamics modeling $\mathcal{L}_{d}(\theta)$ losses with weights $w_{o}$, $w_{r}$, $w_{d}$ to control relative contributions:
$\mathcal{L}(\theta)=w_{o}\mathcal{L}_{o}(\theta)+w_{r}\mathcal{L}_{r}(\theta)+w_{d}\mathcal{L}_{d}(\theta).$ (47)
HarmonyDream (Ma et al., 2024) demonstrated that observation modeling tends to dominate this objective due to visual inputs’ high dimensionality compared to scalar rewards. Their work introduced a variational formulation:
$\mathcal{L}(\theta,w_{o},w_{r},w_{d})$ $=\sum_{i\in\{o,r,d\}}\mathcal{H}(\mathcal{L}_{i}(\theta),\frac{1}{w_{i}})$ (48)
$=\sum_{i\in\{o,r,d\}}w_{i}\mathcal{L}_{i}(\theta)+\log(\frac{1}{w_{i}}),$
where $\mathcal{H}(\mathcal{L}_{i}(\theta),w_{i})=w_{i}\mathcal{L}_{i}(\theta)+\log(1/w_{i})$ dynamically balances the losses by maintaining $\mathbb{E}[w^{*}\cdot\mathcal{L}]=1$. This harmonization technique can substantially enhance sample efficiency and performance. Our work extends these insights through the integration of dynamic task balancing mechanisms into our EDELINE world model architecture.